# Fractionally modulated discrete Carleson’s Theorem and pointwise Ergodic Theorems along certain curves

### v1

@@ 3. Major Box and Multiplier Approximations
The exponential sum estimates allow us to employ a simplified variant of the circle method. The arithmeticity present in the analysis required when dealing with polynomial phases does not enter the picture here. The only frequencies that are expected to be significant for $m^{\mathsf{i}}_{j}$ are the ones close to $(0,0)$. More precisely, let us fix one parameter $\varepsilon\in(0,\min\{1/4,2-c\})$ and define the major box at scale $2^{j}$ as
@@
We collect some useful lemmas to appropriately approximate our multiplier, see Proposition 3.5. In the following four lemmas we have fixed $c\in(1,2)$, $\varepsilon\in(0,\min\{1/4,2-c\})$ and all implicit constants may depend only on these choices.
@@
$m_{j}^{\mathsf{i}}(\xi,\lambda)=\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t){\rm d}t+O\big{(}2^{(2\varepsilon-1)j}\big{)}\text{.}$
@@
We have that
@@
[…]
@@
For every $(\xi,\lambda)\notin\mathfrak{M}_{j}$ and $\mathsf{i}\in\left\{0,1\right\}$, we have that $|m_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim 2^{-\frac{\varepsilon}{4}j}$.
@@
Assume $(\xi,\lambda)\notin\mathfrak{M}_{j}$, by Proposition 2.2 we get
$|m_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim 2^{-j}\min\big{\{}j\big{(}2^{(\frac{c+1}{3})j}+2^{(1-\frac{c}{2})j}|\lambda|^{-\frac{1}{2}}\big{)},|\xi|^{-1}(2^{cj}|\lambda|+1)\big{\}}\text{.}$
@@
$|m_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim j2^{(\frac{c-2}{3})j}+j2^{-\frac{c}{2}j}2^{(-\frac{\varepsilon}{2}+\frac{c}{2})j}\lesssim j2^{(-\frac{2-c}{3})j}+j2^{-\frac{\varepsilon}{2}j}\lesssim 2^{-\frac{\varepsilon}{4}j}\text{,}$
since $\varepsilon<2-c$. On the other hand, if $|\lambda|\leq 2^{(\varepsilon-c)j}$, since $(\xi,\lambda)\notin\mathfrak{M}_{j}$, we get that $|\xi|>2^{(2\varepsilon-1)j}$, and thus
$|m_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim 2^{-j}2^{(1-2\varepsilon)j}(2^{cj}2^{(\varepsilon-c)j}+1)\lesssim 2^{-\varepsilon j}+2^{-2\varepsilon j}\lesssim 2^{-\frac{\varepsilon}{4}j}\text{.}$
@@
(3.1) $\Big{|}\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t){\rm d}t\Big{|}\lesssim 2^{-\frac{\varepsilon}{2}j}\text{.}$
@@
[…]
$\Big{|}\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t){\rm d}t\Big{|}\lesssim|\lambda|^{-\frac{1}{2}}2^{\frac{2-c}{2}j}2^{-j}\lesssim 2^{(-\frac{\varepsilon}{2}+\frac{c}{2})j}2^{-\frac{c}{2}j}\lesssim 2^{-\frac{\varepsilon}{2}j}\text{.}$
@@
$=2^{(2\varepsilon-1)j}(1-c2^{-\varepsilon j})\geq\frac{1}{2}2^{(2\varepsilon-1)j}\text{,}$
since $1-c2^{-\varepsilon j}\geq 1/2\iff j\geq\varepsilon^{-1}\log_{2}(2c)$. We have that $(\lambda t^{c}-\xi t)^{\prime\prime}=\lambda c(c-1)t^{c-2}>0$ and thus the first derivative is monotone, and van der Corput applicable. We get that
@@
$\bigg{|}\int_{\mathbb{R}}e(-\xi t)\psi_{j}(t){\rm d}t\bigg{|}\lesssim|\xi 2^{j}|^{-1}\lesssim 2^{-\frac{\varepsilon}{2}}\text{,}$
and the proof is complete. ∎
@@
Let $\mathsf{i}\in\{0,1\}$. Fix $\eta_{1}\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$ such that $\bm{1}_{[-1/4,1/4]}\leq\eta_{1}\leq\bm{1}_{[-1/2,1/2]}$ and let $\eta(\xi,\lambda)=\eta_{1}(\xi)\eta_{1}(\lambda)$. For every $j\in\mathbb{N}$, let
@@
Then there exists a positive constant $C=C(c,\varepsilon)$ such that
$\sup_{\xi,\lambda\in[-1/2,1/2)}|E_{j}^{\mathsf{i}}(\xi,\lambda)|\leq C2^{-\frac{\varepsilon}{4}j}\quad\text{and}\quad\sup_{\xi,\lambda\in[-1/2,1/2)}|\partial_{\lambda}E_{j}^{\mathsf{i}}(\xi,\lambda)|\leq C2^{cj}\text{.}$
@@
An identical reasoning with the one in the beginning of the proof of Lemma 3.3 shows that it suffices to establish the first estimate for $j\geq 3$ and for all such $j$ we get that $\mathfrak{M}_{j}\subseteq[-1/4,1/4]^{2}$ and thus if $(\xi,\lambda)\in\mathfrak{M}_{j}$, then $\eta(\xi,\lambda)=1$. The first estimate for $(\xi,\lambda)\in\mathfrak{M}_{j}$ is therefore immediate from Lemma 3.1, since for all such $(\xi,\lambda)$ we have $|m^{\mathsf{i}}_{j}(\xi,\lambda)-H^{\mathsf{i}}_{j}(\xi,\lambda)|\lesssim 2^{(2\varepsilon-1)j}\leq 2^{-\frac{\varepsilon}{4}j}$, and to see why the last estimate holds simply note
@@
If $(\xi,\lambda)\notin\mathfrak{M}_{j}$, then Lemmas 3.2,3.3 give that $|m^{\mathsf{i}}_{j}(\xi,\lambda)-H^{\mathsf{i}}_{j}(\xi,\lambda)|\leq|m^{\mathsf{i}}_{j}(\xi,\lambda)|+|H^{\mathsf{i}}_{j}(\xi,\lambda)|\lesssim 2^{-\frac{\varepsilon}{4}j}$. Therefore we have shown that
$\sup_{\xi,\lambda\in[-1/2,1/2)}|E_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim 2^{-\frac{\varepsilon}{4}j}\text{,}$
@@
by its continuous counterpart. More precisely, fix $c\in(1,2)$, $\varepsilon\in(0,\min\{1/4,2-c\})$ and for every $t>1$ let
$\widetilde{\mathfrak{M}}_{t}\coloneqq\big{\{}(\xi_{1},\xi_{2})\in[-1/2,1/2)^{2}:\,|\xi_{1}|\leq t^{2\varepsilon-1},\,|\xi_{2}|\leq t^{\varepsilon-c}\big{\}}\text{.}$
@@
For every $t\in\mathbb{N}$, let $\mathcal{L}_{t}(\xi_{1},\xi_{2}):=\frac{1}{t}\int_{0}^{t}e(-\xi_{2}s^{c}-\xi_{1}s){\rm d}s$. There exists a positive constant $C=C(c,\varepsilon)$ such that
(3.3) $\sup_{\xi_{1},\xi_{2}\in[-1/2,1/2)}|k_{t}(\xi_{1},\xi_{2})-\mathcal{L}_{t}(\xi_{1},\xi_{2})|\leq Ct^{-\frac{\varepsilon}{4}}\text{,}$
@@
(3.4) $\sup_{(\xi_{1},\xi_{2})\in\widetilde{\mathfrak{M}}_{t}}|k_{t}(\xi_{1},\xi_{2})-\mathcal{L}_{t}(\xi_{1},\xi_{2})|\leq Ct^{-\frac{\varepsilon}{4}}\text{,}$
(3.5) $\sup_{(\xi_{1},\xi_{2})\in[-1/2,1/2)^{2}\setminus\widetilde{\mathfrak{M}}_{t}}|k_{t}(\xi_{1},\xi_{2})|\leq Ct^{-\frac{\varepsilon}{4}}\text{,}\quad\text{and}\quad\sup_{(\xi_{1},\xi_{2})\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}}|\mathcal{L}_{t}(\xi_{1},\xi_{2})|\leq Ct^{-\frac{\varepsilon}{4}}\text{.}$
@@
We fix $c$ and $\varepsilon$ as before and all implied constants may depend on them. It suffices to establish the estimates for $t\gtrsim_{c,\varepsilon}1$, and in particular, we will prove the result for $t\geq(c2^{c})^{1/\varepsilon}$. For the first assertion of (3.4) note that for all $(\xi_{1},\xi_{2})\in\widetilde{\mathfrak{M}}_{t}$ we have that
[…]
@@
Now we establish the other two estimates in two steps. Let $\widetilde{\mathfrak{M}}_{t}^{\mathsf{c}}=[-1/2,1/2)\setminus\widetilde{\mathfrak{M}}_{t}$ and let $(\xi_{1},\xi_{2})\in\widetilde{\mathfrak{M}}_{t}^{\mathsf{c}}$. If $|\xi_{2}|>t^{\varepsilon-c}$, then by Proposition 2.1 applied for $N=t$ and $M=\lfloor N^{\frac{2-c}{3}}\rfloor$ we get
$|k_{t}(\xi_{1},\xi_{2})|\lesssim(1+\log t)(t^{\frac{c-2}{3}}+t^{-\frac{c}{2}}t^{\frac{c-\varepsilon}{2}})\lesssim t^{-\frac{\varepsilon}{4}}\text{,}\quad\text{since $\varepsilon<2-c$.}$
If $|\xi_{2}|\leq t^{\varepsilon-c}$, then since $(\xi_{1},\xi_{2})\in\widetilde{\mathfrak{M}}^{\mathsf{c}}_{t}$, we must have that $|\xi_{1}|>t^{2\varepsilon-1}$. But then by the second estimate of Proposition 2.1 we obtain
(3.8) $|k_{t}(\xi_{1},\xi_{2})|\lesssim t^{-1}|\xi_{1}|^{-1}(1+t^{c}|\xi_{2}|)\lesssim t^{-1}t^{1-2\varepsilon}(1+t^{c}t^{\varepsilon-c})\lesssim t^{-2\varepsilon}+t^{-\varepsilon}\lesssim t^{-\frac{\varepsilon}{4}}\text{.}$
Thus we have shown that $\sup_{\xi\in\widetilde{\mathfrak{M}}^{\mathsf{c}}_{t}}|k_{t}(\xi)|\lesssim t^{-\frac{\varepsilon}{4}}$. It remains to show that $\sup_{\xi\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}}|\mathcal{L}_{t}(\xi)|\lesssim t^{-\frac{\varepsilon}{4}}$ and we achieve this via a similar reasoning. Let $(\xi_{1},\xi_{2})\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}$. If $|\xi_{2}|>t^{\varepsilon-c}$, then note that
@@
$\Big{|}\int_{2^{m}}^{\min\{2^{m+1},t\}}e(-\xi_{2}s^{c}-\xi_{1}s){\rm d}s\Big{|}\lesssim|\xi_{2}|^{-1/2}2^{\frac{2-c}{2}m}\lesssim t^{\frac{c-\varepsilon}{2}}2^{\frac{2-c}{2}m}\text{.}$
@@
$|\mathcal{L}_{t}(\xi_{1},\xi_{2})|\lesssim t^{-1}+t^{-1}\sum_{m=0}^{\lfloor\log_{2}t\rfloor}t^{\frac{c-\varepsilon}{2}}2^{\frac{2-c}{2}m}=\lesssim t^{-1}+t^{\frac{c-\varepsilon}{2}-1}\sum_{m=0}^{\lfloor\log_{2}t\rfloor}\big{(}2^{\frac{2-c}{2}}\big{)}^{m}\lesssim t^{-1}+t^{\frac{c-\varepsilon}{2}-1}t^{\frac{2-c}{2}}\lesssim t^{-\frac{\varepsilon}{4}}\text{.}$
If $|\xi_{2}|\leq t^{\varepsilon-c}$, then since $(\xi_{1},\xi_{2})\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}$, we have that $|\xi_{1}|>t^{2\varepsilon-1}$. We will use the first derivative test here. For any $m\in\{0,\dotsc,\lfloor\log_{2}t\rfloor\}$ and $s\in[2^{m},2^{m+1}]$ we have that
$|(-\xi_{2}s^{c}-\xi_{1}s)^{\prime}|\geq|\xi_{1}|-|c\xi_{2}s^{c-1}|>t^{2\varepsilon-1}-ct^{\varepsilon-c}2^{(m+1)(c-1)}\geq\frac{1}{2}t^{2\varepsilon-1}\text{,}$
where for the last estimate we used the fact that $t\geq(c2^{c})^{1/\varepsilon}$. More precisely
$t^{2\varepsilon-1}-ct^{\varepsilon-c}2^{(m+1)(c-1)}\geq t^{2\varepsilon-1}\big{(}1-c2^{c-1}t^{-\varepsilon}\big{)}\geq\frac{1}{2}t^{2\varepsilon-1}\text{,}$
since $1-c2^{c-1}t^{-\varepsilon}\geq 1/2\iff t\geq(c2^{c})^{1/\varepsilon}$. If $\xi_{2}\neq 0$, then the first derivative of the phase is monotonic and by van der Corput we obtain
$|\mathcal{L}_{t}(\xi_{1},\xi_{2})|\lesssim t^{-1}+t^{-1}\sum_{m=0}^{\lfloor\log_{2}t\rfloor}\Big{|}\int_{2^{m}}^{\min\{2^{m+1},t\}}e(-\xi_{2}s^{c}-\xi_{1}s){\rm d}s\Big{|}\lesssim t^{-1}+t^{-1}\sum_{m=0}^{\lfloor\log_{2}t\rfloor}t^{1-2\varepsilon}\lesssim t^{-\frac{\varepsilon}{4}}\text{.}$
@@
[…]
We have shown that $\sup_{\xi\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}}|\mathcal{L}_{t}(\xi)|\lesssim t^{-\frac{\varepsilon}{4}}$ and the proof is complete. ∎

### v2

@@ 3. Major Box and Multiplier Approximations
The exponential sum estimates allow us to employ a simplified variant of the circle method. The arithmeticity present in the analysis required when dealing with polynomial phases does not enter the picture here. The only frequencies that are expected to be significant for $m^{\mathsf{i}}_{j}$ are the ones close to $(0,0)$. More precisely, let us fix one parameter $\varepsilon\coloneqq\frac{2-c}{4}\in(0,\min\{1/4,2-c\})$ [^2] Let us note that if one does not wish to keep track of the dependencies of the strength of the estimates with respect to $c\in(1,2)$, as we do for the majority of the estimates in this section, any choice of $\varepsilon\in(0,\min\{1/4,2-c\})$ will be admissible. and define the major box at scale $2^{j}$ as
@@
We collect some useful lemmas to appropriately approximate our multiplier, see Proposition 3.5.
@@
(3.1) $m_{j}^{\mathsf{i}}(\xi,\lambda)=\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t){\rm d}t+O\big(2^{(2\varepsilon-1)j}\big)\text{.}$
where the implied constant is absolute.
@@
Since we have the trivial estimates $\big|m_{j}^{\mathsf{i}}(\xi,\lambda)\big|\lesssim 1$ and $\Big|\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t){\rm d}t\Big|\lesssim 1$, it suffices to establish (3.1) for $j\geq 3$, and we have
@@
$\Big|m_{j}^{\mathsf{i}}(\xi,\lambda)-\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t)dt\Big|\lesssim\sum_{2^{j-4}\leq|n|\leq 2^{j}}2^{-j}2^{(2\varepsilon-1)j}\lesssim 2^{(2\varepsilon-1)j}\text{,}\quad\text{as desired.}$
@@
For every $j\in\mathbb{N}$, $\mathsf{i}\in\{0,1\}$, and $(\xi,\lambda)\in[-1/2,1/2)^{2}\setminus\mathfrak{M}_{j}$ we have that
$|m_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{4}j}\text{.}$
@@
We can assume that $j\varepsilon\geq 1$ because for $j$ such that $j\varepsilon<1$ the estimate above is immediate from the trivial estimate $|m_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim 1$. Thus Proposition 2.2 is applicable and for $(\xi,\lambda)\notin\mathfrak{M}_{j}$ we get
$|m_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim 2^{-j}\min\big\{(c-1)^{-\frac{1}{2}}(2-c)j\big(2^{(\frac{c+1}{3})j}+2^{(1-\frac{c}{2})j}|\lambda|^{-\frac{1}{2}}\big),|\xi|^{-1}(2^{cj}|\lambda|+1)\big\}\text{.}$
@@
$\frac{|m_{j}^{\mathsf{i}}(\xi,\lambda)|}{(c-1)^{-\frac{1}{2}}}\lesssim(2-c)j2^{(\frac{c-2}{3})j}+(2-c)j2^{-\frac{c}{2}j}2^{(-\frac{\varepsilon}{2}+\frac{c}{2})j}\lesssim\varepsilon j2^{-\frac{\varepsilon}{2}j}\lesssim 2^{-\frac{\varepsilon}{4}j}\text{,}$
where we used that $\varepsilon=\frac{2-c}{4}$, and the fact that
(3.2) $x2^{-\frac{x}{2}}\lesssim 2^{-\frac{x}{4}}\text{.}$
On the other hand, if $|\lambda|\leq 2^{(\varepsilon-c)j}$, since $(\xi,\lambda)\notin\mathfrak{M}_{j}$, we get that $|\xi|>2^{(2\varepsilon-1)j}$, and thus
$|m_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim 2^{-j}2^{(1-2\varepsilon)j}(2^{cj}2^{(\varepsilon-c)j}+1)\lesssim 2^{-\varepsilon j}+2^{-2\varepsilon j}\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{4}j}\text{.}$
@@
(3.3) $\Big|\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t){\rm d}t\Big|\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{2}j}\text{.}$
@@
Using the trivial estimate $\Big|\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t)dt\Big|\lesssim 1$, we see that it suffices to establish (3.3) for $\varepsilon j\gtrsim 1$, in particular we prove it below for $j\geq 2\varepsilon^{-1}$. Taking into account the support of $\psi_{j}$ we may split the integral into the intervals $(-2^{j-1},-2^{j-3})$ and $(2^{j-3},2^{j-1})$ and apply van der Corput’s lemma. Without loss of generality, we may assume that $\lambda\geq 0$. In particular, if $|\lambda|>2^{(\varepsilon-c)j}$ by the second derivative test, see for example Corollary 2.6.8. in [7], one has
$\Big|\int_{\mathbb{R}}e(\lambda[t]_{\mathsf{i}}^{c}-\xi t)\psi_{j}(t){\rm d}t\Big|\lesssim(c(c-1))^{-\frac{1}{2}}|\lambda|^{-\frac{1}{2}}2^{\frac{2-c}{2}j}2^{-j}\lesssim(c-1)^{-\frac{1}{2}}2^{(-\frac{\varepsilon}{2}+\frac{c}{2})j}2^{-\frac{c}{2}j}\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{2}j}\text{.}$
@@
$2^{(2\varepsilon-1)j}(1-c2^{-\varepsilon j})\geq\frac{1}{2}2^{(2\varepsilon-1)j}\text{,}$
since $1-c2^{-\varepsilon j}\geq 1/2\iff j\geq\varepsilon^{-1}\log_{2}(2c)$, which is true because $j\geq 2\varepsilon^{-1}\geq\varepsilon^{-1}\log_{2}(2c)$. We have that $(\lambda t^{c}-\xi t)^{\prime\prime}=\lambda c(c-1)t^{c-2}>0$ and thus the first derivative is monotone, and van der Corput applicable. We get that
@@
$\bigg|\int_{\mathbb{R}}e(-\xi t)\psi_{j}(t){\rm d}t\bigg|\lesssim|\xi 2^{j}|^{-1}\lesssim 2^{-\frac{\varepsilon}{2}j}\text{,}\quad\text{and the proof is complete.}$
@@
Fix $\eta_{1}\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$ such that $\bm{1}_{[-1/4,1/4]}\leq\eta_{1}\leq\bm{1}_{[-1/2,1/2]}$ and let $\eta(\xi,\lambda)=\eta_{1}(\xi)\eta_{1}(\lambda)$. For every $j\in\mathbb{N}$ and $\mathsf{i}\in\{0,1\}$, let
@@
Then we have
$\sup_{\xi,\lambda\in[-1/2,1/2)}|E_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{4}j}\quad\text{and}\quad\sup_{\xi,\lambda\in[-1/2,1/2)}|\partial_{\lambda}E_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim 2^{cj}\text{.}$
@@
A similar reasoning with the one in the beginning of the proof of Lemma 3.3 shows that it suffices to establish the first estimate for $j\geq 3$ and for all such $j$ we get that $\mathfrak{M}_{j}\subseteq[-1/4,1/4]^{2}$ and thus if $(\xi,\lambda)\in\mathfrak{M}_{j}$, then $\eta(\xi,\lambda)=1$. The first estimate for $(\xi,\lambda)\in\mathfrak{M}_{j}$ is therefore immediate from Lemma 3.1, since for all such $(\xi,\lambda)$ we have $|m^{\mathsf{i}}_{j}(\xi,\lambda)-H^{\mathsf{i}}_{j}(\xi,\lambda)|\lesssim 2^{(2\varepsilon-1)j}\leq 2^{-\frac{\varepsilon}{4}j}$, and to see why the last estimate holds simply note
@@
If $(\xi,\lambda)\notin\mathfrak{M}_{j}$, then Lemmas 3.2,3.3 give that
$|m^{\mathsf{i}}_{j}(\xi,\lambda)-H^{\mathsf{i}}_{j}(\xi,\lambda)|\leq|m^{\mathsf{i}}_{j}(\xi,\lambda)|+|H^{\mathsf{i}}_{j}(\xi,\lambda)|\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{4}j}\text{.}$
Therefore we have shown that
$\sup_{\xi,\lambda\in[-1/2,1/2)}|E_{j}^{\mathsf{i}}(\xi,\lambda)|\lesssim(c-1)^{-\frac{1}{2}}2^{-\frac{\varepsilon}{4}j}\text{,}$
@@
by its continuous counterpart. More precisely, fix $c\in(1,2)$, $\tau\in(0,\min\{1/4,2-c\})$ and for every $t>1$ let
$\widetilde{\mathfrak{M}}_{t}\coloneqq\big\{(\xi_{1},\xi_{2})\in[-1/2,1/2)^{2}:\,|\xi_{1}|\leq t^{2\tau-1},\,|\xi_{2}|\leq t^{\tau-c}\big\}\text{.}$
@@
For every $t\in\mathbb{N}$, let $\mathcal{L}_{t}(\xi_{1},\xi_{2}):=\frac{1}{t}\int_{0}^{t}e(-\xi_{2}s^{c}-\xi_{1}s){\rm d}s$. There exists a positive constant $C=C(c,\tau)$ such that
(3.5) $\sup_{\xi_{1},\xi_{2}\in[-1/2,1/2)}|k_{t}(\xi_{1},\xi_{2})-\mathcal{L}_{t}(\xi_{1},\xi_{2})|\leq Ct^{-\frac{\tau}{4}}\text{,}$
@@
(3.6) $\sup_{(\xi_{1},\xi_{2})\in\widetilde{\mathfrak{M}}_{t}}|k_{t}(\xi_{1},\xi_{2})-\mathcal{L}_{t}(\xi_{1},\xi_{2})|\leq Ct^{-\frac{\tau}{4}}\text{,}$
(3.7) $\sup_{(\xi_{1},\xi_{2})\in[-1/2,1/2)^{2}\setminus\widetilde{\mathfrak{M}}_{t}}|k_{t}(\xi_{1},\xi_{2})|\leq Ct^{-\frac{\tau}{4}}\text{,}\quad\text{and}\quad\sup_{(\xi_{1},\xi_{2})\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}}|\mathcal{L}_{t}(\xi_{1},\xi_{2})|\leq Ct^{-\frac{\tau}{4}}\text{.}$
@@
We fix $c$ and $\tau$ as before and all implied constants here may depend on them. It suffices to establish the estimates for $t\gtrsim_{c,\tau}1$, and in particular, we will prove the result for $t\geq(c2^{c})^{1/\tau}$. For the first assertion of (3.6) note that for all $(\xi_{1},\xi_{2})\in\widetilde{\mathfrak{M}}_{t}$ we have that
(3.8) $t|k_{t}(\xi_{1},\xi_{2})-\mathcal{L}_{t}(\xi_{1},\xi_{2})|=\bigg|\sum_{m=1}^{t}\int_{m-1}^{m}\big(e(-\xi_{1}m-\xi_{2}\lfloor m^{c}\rfloor)-e(-\xi_{1}s-\xi_{2}s^{c})\big){\rm d}s\bigg|\\ \leq\sum_{m=1}^{t}\int_{m-1}^{m}\big|e(-\xi_{1}m-\xi_{2}\lfloor m^{c}\rfloor)-e(-\xi_{1}s-\xi_{2}s^{c})\big|{\rm d}s\\ \lesssim\sum_{m=1}^{t}\int_{m-1}^{m}\Big(|\xi_{1}|(m-s)+|\xi_{2}|(\lfloor m^{c}\rfloor-s^{c})\Big)ds\lesssim\sum_{m=1}^{t}\big(t^{2\tau-1}+t^{\tau-c}m^{c-1}\big)\lesssim t^{2\tau}+t^{\tau}\lesssim t^{2\tau}\text{.}$
@@
Now we establish the other two estimates in two steps. Let $\widetilde{\mathfrak{M}}_{t}^{\mathsf{c}}=[-1/2,1/2)\setminus\widetilde{\mathfrak{M}}_{t}$ and let $(\xi_{1},\xi_{2})\in\widetilde{\mathfrak{M}}_{t}^{\mathsf{c}}$. If $|\xi_{2}|>t^{\tau-c}$, then by Proposition 2.1 applied for $N=t$ and $M=\lfloor N^{\frac{2-c}{3}}\rfloor$ we get
$|k_{t}(\xi_{1},\xi_{2})|\lesssim(1+\log t)(t^{\frac{c-2}{3}}+t^{-\frac{c}{2}}t^{\frac{c-\tau}{2}})\lesssim t^{-\frac{\tau}{4}}\text{,}\quad\text{since $\tau<2-c$.}$
If $|\xi_{2}|\leq t^{\tau-c}$, then since $(\xi_{1},\xi_{2})\in\widetilde{\mathfrak{M}}^{\mathsf{c}}_{t}$, we must have that $|\xi_{1}|>t^{2\tau-1}$. But then by the second estimate of Proposition 2.1 we obtain
(3.10) $|k_{t}(\xi_{1},\xi_{2})|\lesssim t^{-1}|\xi_{1}|^{-1}(1+t^{c}|\xi_{2}|)\lesssim t^{-1}t^{1-2\tau}(1+t^{c}t^{\tau-c})\lesssim t^{-2\tau}+t^{-\tau}\lesssim t^{-\frac{\tau}{4}}\text{.}$
Thus we have shown that $\sup_{\xi\in\widetilde{\mathfrak{M}}^{\mathsf{c}}_{t}}|k_{t}(\xi)|\lesssim t^{-\frac{\tau}{4}}$. It remains to show that $\sup_{\xi\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}}|\mathcal{L}_{t}(\xi)|\lesssim t^{-\frac{\tau}{4}}$ and we achieve this via a similar reasoning. Let $(\xi_{1},\xi_{2})\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}$. If $|\xi_{2}|>t^{\tau-c}$, then note that
@@
$\Big|\int_{2^{m}}^{\min\{2^{m+1},t\}}e(-\xi_{2}s^{c}-\xi_{1}s){\rm d}s\Big|\lesssim|\xi_{2}|^{-1/2}2^{\frac{2-c}{2}m}\lesssim t^{\frac{c-\tau}{2}}2^{\frac{2-c}{2}m}\text{.}$
@@
$|\mathcal{L}_{t}(\xi_{1},\xi_{2})|\lesssim t^{-1}+t^{-1}\sum_{m=0}^{\lfloor\log_{2}t\rfloor}t^{\frac{c-\tau}{2}}2^{\frac{2-c}{2}m}=\lesssim t^{-1}+t^{\frac{c-\tau}{2}-1}\sum_{m=0}^{\lfloor\log_{2}t\rfloor}\big(2^{\frac{2-c}{2}}\big)^{m}\lesssim t^{-1}+t^{\frac{c-\tau}{2}-1}t^{\frac{2-c}{2}}\lesssim t^{-\frac{\tau}{4}}\text{.}$
If $|\xi_{2}|\leq t^{\tau-c}$, then since $(\xi_{1},\xi_{2})\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}$, we have that $|\xi_{1}|>t^{2\tau-1}$. We will use the first derivative test here. For any $m\in\{0,\dotsc,\lfloor\log_{2}t\rfloor\}$ and $s\in[2^{m},2^{m+1}]$ we have that
$|(-\xi_{2}s^{c}-\xi_{1}s)^{\prime}|\geq|\xi_{1}|-|c\xi_{2}s^{c-1}|>t^{2\tau-1}-ct^{\tau-c}2^{(m+1)(c-1)}\geq\frac{1}{2}t^{2\tau-1}\text{,}$
where for the last estimate we used the fact that $t\geq(c2^{c})^{1/\tau}$. More precisely
$t^{2\tau-1}-ct^{\tau-c}2^{(m+1)(c-1)}\geq t^{2\tau-1}\big(1-c2^{c-1}t^{-\tau}\big)\geq\frac{1}{2}t^{2\tau-1}\text{,}$
since $1-c2^{c-1}t^{-\tau}\geq 1/2\iff t\geq(c2^{c})^{1/\tau}$. If $\xi_{2}\neq 0$, then the first derivative of the phase is monotonic and by van der Corput we obtain
$|\mathcal{L}_{t}(\xi_{1},\xi_{2})|\lesssim t^{-1}+t^{-1}\sum_{m=0}^{\lfloor\log_{2}t\rfloor}\Big|\int_{2^{m}}^{\min\{2^{m+1},t\}}e(-\xi_{2}s^{c}-\xi_{1}s){\rm d}s\Big|\lesssim t^{-1}+t^{-1}\sum_{m=0}^{\lfloor\log_{2}t\rfloor}t^{1-2\tau}\lesssim t^{-\frac{\tau}{4}}\text{.}$
@@
(3.12) $|\mathcal{L}_{t}(\xi_{1},\xi_{2})|=t^{-1}\Big|\int_{0}^{t}e(-\xi_{1}s)ds\Big|\lesssim t^{-1}|\xi_{1}|^{-1}\lesssim t^{-2\tau}\lesssim t^{-\frac{\tau}{4}}\text{.}$
We have shown that $\sup_{\xi\in\mathbb{R}^{2}\setminus\widetilde{\mathfrak{M}}_{t}}|\mathcal{L}_{t}(\xi)|\lesssim t^{-\frac{\tau}{4}}$ and the proof is complete. ∎